### v1

# L-PR: Exploiting LiDAR Fiducial Marker for Unordered Low Overlap Multiview Point Cloud Registration

Abstract
Point cloud registration is a prerequisite for many applications in computer vision and robotics. Most existing methods focus on pairwise registration of two point clouds with high overlap. Although there have been some methods for low overlap cases, they struggle in degraded scenarios. This paper introduces a novel framework named L-PR, designed to register unordered low overlap multiview point clouds leveraging LiDAR fiducial markers. We refer to them as LiDAR fiducial markers, but they are the same as the popular AprilTag and ArUco markers—thin sheets of paper that do not affect the 3D geometry of the environment. We first propose an improved adaptive threshold marker detection method to provide robust detection results when the viewpoints among point clouds change dramatically. Then, we formulate the unordered multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop a framework consisting of two levels of graphs to address it. The first-level graph, constructed as a weighted graph, is designed to efficiently and optimally infer initial values of scan poses from the unordered set. The second-level graph is constructed as a factor graph. By globally optimizing the variables on the graph, including scan poses, marker poses, and marker corner positions, we tackle the MAP problem. We conduct qualitative and quantitative experiments to demonstrate that the proposed method exhibits superiority over competitors in four aspects: (1) registration accuracy, (2) instance reconstruction quality, (3) localization accuracy, and (4) robustness to the degraded scene. To benefit the community, we open-source our method and dataset at https://github.com/yorklyb/LiDAR-SFM.
Index Terms:
LiDAR Fiducial Marker; Multiview Point Cloud; Low Overlap; Registration.
I Introduction
Point cloud registration is a fundamental need in computer vision [1, 2] and robotics [3, 4, 5]. Most existing point registration methods [3, 6, 7] focus on pairwise registration of two point clouds with high overlap, which are not robust in practical applications [2]. Some recent learning-based research proposes multiview registration methods [1, 8] and studies low overlap cases [2, 1]. Despite their promising performance on benchmarks, the generalization of learning-based methods to unseen scenarios (i.e., out-of-distribution cases of training data) remains problematic. Moreover, both geometry-based methods [3, 4, 5, 9] and learning-based methods [1, 2, 7] rely on geometric features in the scenes to align point clouds. Hence, these methods struggle in degraded scenes [10], such as those with repetitive structures and weak textures.
Figure 1: (a): Overview of L-PR. (1) Given an unordered set of unaligned, low overlap 3D point clouds, we aim to register them into (5) a complete point cloud utilizing LiDAR fiducial markers (thin sheets of paper/board attached to other planes). This is achieved through the proposed (2) adaptive threshold marker detection method and the (3)(4) two-level graphs. The (3) first-level graph handles the unordered set of point clouds and calculates initial values. The (4) second-level graph registers the point clouds by finding the optimal solution to a maximum a-posteriori problem. (b): The proposed method is robust to degraded scenes, such as scenes with repetitive structures.
Reviewing the domain of 2D images, it is found that fiducial markers, such as AprilTag [11] and ArUco [12], have been widely used to tackle challenging situations, such as dramatic viewpoint changes (low overlap) between frames and textureless environments [13, 14, 15]. The visual fiducial markers have rich patterns that are robust for detection and matching, thanks to elaborately designed encoding-decoding algorithms [11, 12]. They pose as thin sheets of paper that can be attached to surfaces without disrupting the 3D environment. In contrast, fiducial objects developed for LiDARs, such as calibration boards [16, 17, 18], lack encoding-decoding pattern design and often take the form of large boards mounted on tripods, thus affecting the 3D environment. LiDARTag [19] is the first work to develop a fiducial marker system akin to AprilTag [11] for LiDARs. While its patterns are compatible with AprilTag, the marker itself is still an additional object placed on a tripod. Recently, the intensity image-based LiDAR fiducial marker system (IFM) [20] proposes to utilize intensity and range images to detect markers. Thus, IFM is compatible with popular visual fiducial markers [11, 12] and inherits the virtue of not affecting the 3D environment. However, the marker detection of IFM is not robust when the viewpoints change in a wide range [20], which hinders its application to in-the-wild multiview point cloud registration. Although the development of LiDAR fiducial markers [19, 20] has advanced, their applications are limited to calibration [21] and AR [20]. The utilization of LiDAR fiducial markers in multiview point cloud registration remains an open problem.
In this paper, we develop a novel framework, as shown in Fig. 1(a), to exploit LiDAR fiducial markers for unordered low overlap multiview point cloud registration. As seen, the LiDAR fiducial markers are thin sheets of board/paper attached to other planes without an impact on the environment. Given an unordered set of low overlap point clouds, the proposed method, dubbed L-PR, efficiently registers them in a global frame. To address the unstable marker detection of IFM [20] in the wild, we first develop an adaptive threshold marker detection method that is robust to viewpoint switching. Then, we formulate the multiview point cloud registration problem as a maximum a-posteriori (MAP) problem and develop two-level graphs to tackle it. The first-level graph, constructed as a weighted graph, is designed to efficiently process the unordered set of point clouds and calculate the initial values of poses among point clouds. In particular, the weights represent the pose estimation error of each marker observation, and we obtain the optimal solutions of initial values by finding the shortest path from an anchor scan to each non-anchor scan on the graph. Given the initial values, we design the second-level graph, a factor graph, to resolve the MAP problem by globally optimizing the poses of point clouds, markers, and positions of marker corners. We conduct both qualitative and quantitative experiments to demonstrate the superiority of the proposed method over competitors [1, 3, 22, 5, 23, 4, 15] in four aspects: (1) registration accuracy, (2) instance reconstruction quality, (3) localization accuracy, and (4) robustness to the degraded scene (see Fig. 1(b)).
The contributions of this work are threefold:
  • We design a novel framework, named L-PR, for unordered low overlap multiview point cloud registration leveraging LiDAR fiducial markers.
  • We develop an improved LiDAR fiducial marker detection algorithm which is robust to viewpoint changes.
  • We release the code implementation and datasets at https://github.com/yorklyb/LiDAR-SFM.
II Related Works
LiDAR Fiducial Markers. Given the success of Structure from Motion [13, 15] and visual Simultaneous Localization and Mapping (SLAM) [14, 24] approaches based on visual fiducial markers [11, 12], why do we still need to exploit LiDAR and LiDAR fiducial markers? There are three reasons. (1) Unlike image-based fiducial marker detection, which is sensitive to ambient light, LiDAR fiducial marker detection is robust to unideal illumination conditions [19, 20], such as purely dark or overexposed environments. (2) Unlike image-based fiducial marker detection, which suffers from the rotational ambiguity problem [13, 15, 14], LiDAR fiducial marker pose estimation is not affected by rotational ambiguity [19, 20]. (3) Data captured by LiDARs is irreplaceable by data acquired from cameras due to the unique modality of LiDAR point cloud [25].
Figure 2: (a) Typical calibration board: forms a thick board with holes and/or regions with high-intensity materials. It impacts the 3D environment and does not support patterns generated by an encoding-decoding algorithm. (b) AprilTag [11]: has rich patterns generated from the encoding-decoding algorithm and is thus robust to false positives/negatives. It has no effect on the 3D environment but is designed for cameras. (c) LiDARTag [19]: has patterns compatible with AprilTag [11] but still forms an additional 3D object. (d) IFM [20]: has patterns compatible with AprilTag [11] and ArUco [12]. It has no impact on the 3D environment.
Most LiDAR fiducial objects [16, 17, 18] are designed for calibration purposes. Fig. 2(a) shows a typical calibration board, a thick board with holes and/or regions covered by high-intensity materials. A calibration board is commonly placed on a tripod for spatial distinction, making it an additional 3D object that affects the environment. Moreover, while a calibration board can provide fiducials (holes, corners, and high-intensity regions) and it is feasible to assign specific indexes to the fiducials, a calibration board is fundamentally different from fiducial marker systems [11, 12]. The core of fiducial marker systems [11, 12] lies in the elaborately designed encoding-decoding algorithms, leveraging which various complex patterns are generated as depicted in Fig. 2(b). In addition to supporting a larger number of unique IDs, the patterns of fiducial markers are robust against false positive/negative issues. Recently, utilizing the encoding-decoding algorithm of AprilTag 3 [11], the first fiducial marker system for LiDARs, named LiDARTag, is introduced in [19]. Unfortunately, as presented in Fig. 2(c), a LiDARTag is still an additional 3D object impacting the 3D environment. IFM [20] develops a LiDAR fiducial marker system that can be integrated with various visual fiducial marker systems [11, 12]. Furthermore, as seen in Fig. 2(d), the usage/placement of IFM is as convenient as the visual fiducial marker. Nevertheless, the marker detection of IFM is unstable when the viewpoints change dramatically among scans due to the use of a constant threshold to process the intensity image [20]. We develop an improved adaptive threshold marker detection method to address this problem because robust marker detection is the prerequisite for exploiting markers in point cloud registration.
Point Cloud Registration. Geometry-based point cloud registration methods, such as variants of iterative closest point methods [26, 4, 5], RANSAC [9], and Teaser++ [3], mainly focus on designing efficient and robust algorithms for describing geometry and extracting geometric features (e.g., points, lines, and surfaces/normals) to find correspondence between point clouds and align them. Recently, Traj LO [22] proposes to utilize temporal information in addition to geometric information to boost the registration of real-time LiDAR scans. In learning-based point cloud registration methods like Predator [2] and SGHR [1], neural networks are designed to learn features representing the geometry [7, 27, 8, 6], and then learn to utilize these features for registering the point clouds. While these methods demonstrate remarkable performance on benchmarks, they still face challenges in handling degraded scenes, such as repetitive structures and textureless environments. Moreover, most existing methods [26, 4, 5, 7, 6] only apply to high overlap scenarios. Despite the existence of some research [1, 2] on low overlap cases, their generalization to unseen scenes is limited. Inspired by the utilization of visual fiducial markers in camera-based applications [13, 15, 14, 24] to handle degraded scenes and low overlap cases, we propose exploiting LiDAR fiducial markers to enhance multiview point cloud registration.
III Methodology
III-A Adaptive Threshold LiDAR Fiducial Marker Detection
Preliminaries. Given a 3D point cloud, IFM [20] generates a 2D intensity image by mapping each 3D point to a pixel on an image plane:
$u=\lceil\frac{\theta}{\alpha_{a}}\rfloor+u_{o},\;v=\lceil\frac{\phi}{\alpha_{i}}\rfloor+v_{o},$ (1)
where $[u,v]^{T}$ denote the image coordinates of the projected pixel. $[\theta,\phi,r]^{T}$ denote the spherical coordinates of the 3D point. $\alpha_{a}$ and $\alpha_{i}$ represent the angular resolutions in $u$ (azimuth) and $v$ (inclination) directions, respectively. $u_{o}$ and $v_{o}$ are the offsets. The pixel value $\mathbf{I}(u,v)$ is decided by the corresponding intensity value through the color map. The ranging information is also saved for each pixel. In the real world, the intensity values of the same surface of an object could vary, leading to noise in the rendered image. To denoise the intensity image, each pixel is binarized with a threshold $\lambda$. In [20], $\lambda$ is set as a constant. Then, the embedded 2D marker detector [11, 12] is applied to localize the fiducials on the binarized image. Using the ranging information, the detected 2D fiducials are projected back into 3D space via the reverse of Eq. (1) and become 3D fiducials.
Adaptive Threshold Method. As aforementioned, binarization is applied to the raw intensity image due to the imaging noise (see the zoomed views of Fig. 3). As shown in Fig. 3, the effect is determined by the threshold $\lambda$. In [20], it is found that $\lambda$ can be selected as a constant if the viewpoint does not change drastically. However, with significant changes in the scene, the value of $\lambda$ needs adjustment. This poses a challenge for multiview point cloud registration as the viewpoints of scans can vary dramatically. Therefore, unless an appropriate $\lambda$ is carefully selected for each scan, the LiDAR fiducial marker detection will fail. Take the case shown in Fig. 3 as an example, $\lambda$=12.5 demonstrates the best denoising performance, while other values result in image degradation. Thus, $\lambda$=12.5 is the optimal threshold compared to other values, denoted by $\lambda^{*}$. Unlike [20], which determines $\lambda^{*}$ based on experience or experimentation, we develop an algorithm to automatically seek $\lambda^{*}$.
Figure 3: The raw intensity image binarized with different threshold values.
Given that our focus is on marker detection rather than image denoising, we design Algorithm 1 that utilizes the detection result as the feedback to search for $\lambda^{*}$ automatically. In particular, we design a loop, in which we gradually increase $\lambda$, to determine the threshold value that maximizes the number of detected markers. We experimentally found that it is more efficient to search for $\lambda$ starting from a small value (see Fig. 3). After applying Algorithm 1, the 2D fiducials located on the image binarized with $\lambda^{*}$ are projected into 3D fiducials using the subsequent steps of IFM [20]. Our algorithm resolves the problem of a constant threshold being inapplicable to all scans, especially when LiDAR moves in the wild and scenes undergo significant changes.
Input: Raw intensity image, $\mathbf{I}$
Output: The optimal threshold, $\lambda^{*}$.
Initialize parameters: Search scope, $S$. step size, $\delta$. queue for saving detected markers, $Q=[\ ]$. length of $Q$, $Q_{l}=0$. temporary queue, $Q_{temp}=[\ ]$. length of $Q_{temp}=[\ ]$, $Q_{temp,l}=0$. the optimal threshold, $\lambda^{*}=0$. search step, $i=0$.
Define the binarization operation as $\Psi(\mathbf{I},\lambda)$.
Define the marker detector operation as $\Gamma(\mathbf{I})$.
while $i<S$ do
       $\lambda=\delta\times i$
       $\mathbf{I}=\Psi(\mathbf{I},\lambda)$
       $\Gamma(\mathbf{I})\rightarrow Q_{temp}$
       if $Q_{temp,l}\geq Q_{l}$ then
             for marker in $Q_{temp}$ do
                   if marker not in $Q$ then
                        append marker to $Q$
            $\lambda^{*}=\lambda$
      $i=i+1$
Return the image binarized with $\lambda^{*}$.
Algorithm 1 Search for the optimal threshold, $\lambda^{*}$.
III-B Problem Formulation
In this section, we introduce the problem formulation. Suppose that the marker size is $l$. In the marker coordinate system $\{M\}_{j}$, we know that the 3D coordinates of the four corners of the $j$-th marker, ${}^{j}\mathbf{p}^{j,1},^{j}\mathbf{p}^{j,2},^{j}\mathbf{p}^{j,3},$ and ${}^{j}\mathbf{p}^{j,4}$, are $[-l/2,-l/2,0]^{T},[l/2,-l/2,0]^{T},[l/2,l/2,0]^{T},$ and $[-l/2,l/2,0]^{T}$, respectively. LiDAR fiducial marker detection returns the 3D coordinates of the corners expressed in $\{F\}_{i}$ (the local coordinate system of the $i$-th scan $f_{i}$) [20]. Thus, the 6-DOF transmission from $\{M\}_{j}$ to $\{F\}_{i}$, denoted as $\mathbf{T}^{j}_{i}$, can be found by solving the following least square problem:
$\mathbf{T}^{j,*}_{i}=\underset{\mathbf{T}^{j}_{i}}{\arg\min}\sum_{s=1}^{4}\left\|\mathbf{T}^{j}_{i}\ \cdot\ ^{j}\mathbf{p}^{j,s}-_{i}\mathbf{p}^{j,s}\right\|^{2}.$ (2)
The Singular Value Decomposition (SVD) method [28] is employed to compute $\mathbf{T}^{j,*}_{i}$. We denote the set of measurements, including (1) the corner positions w.r.t. $\{M\}_{j}$, (2) the corner positions w.r.t. $\{F\}_{i}$, and (3) the marker poses w.r.t. $\{F\}_{i}$, as $\mathcal{Z}$.
Figure 4: An illustration of the first-level graph. In this simple case, as shown on the right side, four alternative paths exist from $f_{2}$ to the anchor scan ($f_{1}$). We employ Dijkstra’s algorithm [29] to find the shortest path.
To register the multiview point clouds, we need to find a globally consistent pose for every scan so that the point clouds can be transformed into a complete point cloud in the global coordinate system $\{G\}$. To achieve this objective, we consider the following variables in this work: (1) the poses of point clouds w.r.t. $\{G\}$, (2) the poses of markers w.r.t. $\{G\}$, and (3) the marker corner positions w.r.t. $\{G\}$. We specify the set of variables as $\Theta$. Finally, a maximum a-posteriori (MAP) inference problem is formulated: given the measurements. $\mathcal{Z}$, the goal is to find the optimal variable set $\Theta^{*}$ that maximizes the posterior probability $P(\Theta\mid\mathcal{Z})$:
$\ \Theta^{*}=\underset{\Theta}{\arg\max}\ P(\Theta\mid\mathcal{Z}).$ (3)
In the following, we design a framework consisting of two-level graphs to resolve this MAP problem.
III-C First-level Graph
As aforementioned, the variables in $\Theta$ to be optimized cannot be directly obtained through the measurements in $\mathcal{Z}$. Moreover, deriving the initial values of the variables requires both efficiency and accuracy, given that the input is an unordered set of low overlap point clouds. To address this challenge, we design the first-level graph. First, we consider computing the relative pose between two point clouds. Suppose the $j$-th marker appears in the scenes of two scans $f_{i}$ and $f_{m}$, $\mathbf{T}^{j}_{i}$ and $\mathbf{T}^{j}_{m}$ can be calculated using Eq. (2). Consequently, the relative pose between $f_{i}$ and $f_{m}$, denoted as $\mathbf{T}_{i,m}$, is available from
$\mathbf{T}_{i,m}=(\mathbf{T}^{j}_{i})^{-1}\mathbf{T}^{j}_{m},$ (4)
where $(\mathbf{T}^{j}_{i})^{-1}$ indicates the inverse of $\mathbf{T}^{j}_{i}$. $(\mathbf{T}^{j}_{i})^{-1}$ indicates the pose that transforms 3D points from $\{F\}_{i}$ to $\{M\}_{j}$. Although the method introduced in Eq. (4) is straightforward, it cannot be directly applied because the input in this work is an unordered set of point clouds that do not follow a temporal or spatial sequence. Specifically, for scans with no shared marker observations, their relative pose has to be calculated through pose propagation among other scans. However, multiple alternative paths could exist. Even for scans that share marker observations, there could be more than one overlapped marker. Hence, to accurately and efficiently estimate relative poses among scans, it is necessary to design an algorithm to determine which scans and markers to apply Eq. (4).
Specifically, our objective is to infer the relative poses with the highest possible accuracy using only the necessary low-dimensional information through a simple process. Thus, we construct the first-level graph as a weighted graph. The construction of this graph is as follows. When processing the point clouds with the proposed adaptive threshold method one by one, if a marker is detected in a point cloud, we add the corresponding scan node and marker node to the graph along with a weighted edge. The edge weight is the pose estimation point-to-point error ${e}_{pp}$:
${e}_{pp}=\sum_{s=1}^{4}\left\|\mathbf{T}^{j,*}_{i}\ \cdot\ ^{j}\mathbf{p}^{j,s}-_{i}\mathbf{p}^{j,s}\right\|^{2}.$ (5)
Eq. (5) indicates the substitution of the result given by SVD back into the right side of Eq. (2). ${e}_{pp}\in\mathbb{R}^{+}$ is employed as the metric to evaluate the quality of pose estimation of the marker in the corresponding point cloud. We define the first scan as the anchor scan. The local coordinate system of the anchor scan is set as the global coordinate system. Namely, $\{F\}_{1}=\{G\}$. We only consider the relative poses between the anchor scan and each non-anchor scan. Although there could be multiple paths from the given scan to the anchor scan (see Fig. 4), we have already saved the information on pose estimation quality by constructing the first-level graph as a weighted graph. Therefore, we adopt Dijkstra’s algorithm [29] to obtain the shortest path. Then, we compute the relative pose along the shortest path iteratively using Eq. (4). Given that the relative pose computation is achieved with the lowest accumulation of $e_{pp}$, the estimation result achieves the highest possible accuracy. Moreover, the search for the optimal path to propagate poses is based merely on the one-dimensional $e_{pp}$, without incorporating any 6-DOF poses or 3D locations.
In this way, the point cloud poses w.r.t. $\{G\}$ are obtained. Since the marker detection provides the marker poses and the 3D positions of marker corners w.r.t. the local coordinate system of corresponding scan, we can derive the initial values of the marker poses w.r.t. $\{G\}$ and the corner positions w.r.t. $\{G\}$ through the point cloud poses w.r.t. $\{G\}$.
TABLE I: quantitative comparison with SGHR [1] and Teaser++ [3] on our test dataset
Scene ##\## 1 2 3 4 5 6 7 8 9 10
Teaser++[3] $\mathrm{RMSE}_{R}$ (rad) 1.504 1.795 2.394 2.277 2.493 1.946 1.632 1.818 1.779 1.876
$\mathrm{RMSE}_{T}$ (m) 1.724 1.890 2.075 2.252 1.911 1.733 2.089 1.861 1.848 1.891
Time (s) 313.1 298.3 336.4 322.3 388.9 419.8 441.5 397.1 327.2 401.2
SGHR[1] $\mathrm{RMSE}_{R}$ (rad) 0.152 0.060 1.198 1.825 1.311 1.696 4.42 2.792 3.034 2.810
$\mathrm{RMSE}_{T}$ (m) 0.020 0.010 0.094 0.156 0.231 0.274 0.949 2.171 0.686 1.787
Time (s) 846.1 785.0 886.5 825.3 851.7 979.3 983.6 909.9 794.6 950.4
Ours $\mathrm{RMSE}_{R}$ (rad) 0.036 0.068 0.089 0.065 0.088 0.078 0.067 0.069 0.087 0.101
$\mathrm{RMSE}_{T}$ (m) 0.017 0.011 0.028 0.031 0.032 0.043 0.019 0.082 0.069 0.077
Time (s) 31.1 21.2 35.7 32.6 36.2 43.5 53.8 39.9 24.9 42.2
Figure 5: The procedures for formulating the second graph. The variable nodes are represented by circles and the factor nodes are represented by squares.
III-D Second-level Graph
In this section, we construct a factor graph to optimize the variables, ultimately addressing the MAP inference problem introduced in Eq. (3). Suppose that the $j$-th marker is detected in the $i$-th scan, the following nodes are added to the second graph. The variable nodes include $\mathbf{T}^{j}$ (6-DOF pose of the $j$-th marker w.r.t. $\{G\}$), $\mathbf{T}_{i}$ (6-DOF pose of the $i$-th scan w.r.t. $\{G\}$), and $\{\mathbf{p}^{j,1},\mathbf{p}^{j,2},\mathbf{p}^{j,3},\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{G\}$). The factor nodes include $\mathbf{T}^{j}_{i}$ (6-DOF pose of the $j$-th marker w.r.t. $\{F\}_{i}$), $\{^{j}\mathbf{p}^{j,1},\ ^{j}\mathbf{p}^{j,2},\ ^{j}\mathbf{p}^{j,3},\ ^{j}\mathbf{p}^{j,4}\}$ (3D coordinates of the coroners of the $j$-th marker w.r.t. $\{M\}_{j}$), and $\{_{i}\mathbf{p}^{j,1},\ _{i}\mathbf{p}^{j,2},$ ${}_{i}\mathbf{p}^{j,3},$ ${}_{i}\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{F\}_{i}$).
For clarity, the first stage in Fig. 5 shows the added nodes and edges when a marker is detected in a scan. The graph becomes the one depicted in the second stage of Fig. 5 after integrating all the marker detection results. Subsequently, we add a prior factor that connects to the first (anchor) point cloud pose node since we set $\{F\}_{1}$ as $\{G\}$. Finally, we add factor nodes representing the relative poses between the anchor scan and each non-anchor scan. Up to this point, the factor graph is completed, as shown in stage 3 of Fig. 5. Following [30, 31], this indicates that the factorization of the posterior probability $P(\Theta\mid\mathcal{Z})$ has been defined:
$P(\Theta\mid\mathcal{Z})=\prod_{k}P^{(k)}(\Theta),$ (6)
where $P^{(k)}$ are the factors in the second graph and we follow the standard pipeline [30] to model them as Gaussians:
$P^{(k)}(\Theta)\propto\exp\left(-\frac{1}{2}\left\|h_{k}\left(\Theta\right)\ominus z_{k}\right\|_{\Sigma_{k}}^{2}\right),$ (7)
where $h_{k}\left(\Theta\right)$ is the measurement function and $z_{k}$ is a measurement. $\|e\|_{\Sigma}^{2}\triangleq e^{T}\Sigma^{-1}e$ denotes the squared Mahalanobis distance with $\Sigma$ being the covariance matrix. Following [24], if $z_{k}\in\mathbb{R}^{3\times 1}$ is a 3D position, $\ominus$ refers to straight subtraction for elements. If $z_{k}\in SE(3)$ is a 6-DOF pose, $\ominus$ generates 6-dimensional Lie algebra coordinates:
$\mathbf{T}_{A}\ominus\mathbf{T}_{B}=[[\log(\mathrm{Rot}(\mathbf{T}_{B}^{-1}\mathbf{T}_{A}))]^{T}_{\vee},\mathrm{Trans}(\mathbf{T}_{B}^{-1}\mathbf{T}_{A})^{T}]^{T},$
where for a $\mathbf{T}\in SE(3)$, $\mathrm{Rot}(\mathbf{T})$ denotes the rotation matrix $\mathbf{R}\in SO(3)$ and $\mathrm{Trans}(\mathbf{T})$ denotes translation vector $\mathbf{t\in\mathbb{R}^{3\times 1}}$. $\log(\cdot)$ represents the matrix logarithm. $\vee$ is the vee map operator that finds the unique vector $\xi\in\mathbb{R}^{3\times 1}$ (Lie algebra coordinates) corresponding to a given skew-symmetric matrix $\log(\mathbf{R})\in\mathbb{R}^{3\times 3}$ [28, 24]. With the Gaussian modeling, the objective function in Eq. (6) is transformed into a least square problem by applying the negative logarithm:
$\arg\min_{\Theta}(-\log\prod_{k}P^{(k)}(\Theta))=\arg\min_{\Theta}\frac{1}{2}\sum_{k}\left\|h_{k}\left(\Theta\right)\ominus z_{k}\right\|_{\Sigma_{k}}^{2}.$
Figure 6: An illustration of $\lambda^{*}$ across various scenes and positions.
We utilize the Levenberg-Marquardt algorithm [31] to resolve this problem. The acquisition of initial values is introduced in Section III-C. The noise covariance matrices are determined by the quantitative experiments conducted in [20].
Remark. To clarify, the difference between IFM [20] and this work resembles the difference between AprilTag [11] and marker-based SfM [15, 13]. Although IFM and AprilTag address the problem of sensor pose estimation, constructing an efficient framework for multiview point cloud registration or SfM based on markers is a different endeavor, which is the contribution of [15, 13] as well as this work.
IV Experimental Results
In this section, we report the experimental results of five tests to demonstrate the effectiveness of the proposed method from different perspectives.
IV-A Evaluation of the Adaptive Threshold Method
Data. We place the Livox MID-40 LiDAR and a 69.2 cm $\times$ 69.2 cm ArUco marker in three different scenes as shown in Fig. 6(a). The three scenes are between groups of buildings, open outdoors, and large indoor parking lots, respectively. In each scene, we collect point clouds and test the proposed adaptive threshold method at different relative positions. Results and Analysis. The 3D stem plots in Fig. 7 present the results, where x and y indicate the relative position of the LiDAR in the marker coordinate system, and the z-axis corresponds to $\lambda^{*}$. In scene 1, we test at 21 different locations, and $\lambda^{*}$ varies from 7 to 36. In scene 2, we test at 21 viewpoints, and $\lambda^{*}$ varies from 6 to 15. In scene 3, we test at 9 positions, and $\lambda^{*}$ varies from 6 to 20. The experiments carried out in this section demonstrate that when the viewpoint changes dramatically, a constant threshold [20] is not applicable. Thus, the proposed adaptive threshold method is necessary.
Figure 7: Visual comparison with SOTAs (SGHR [1], Teaser++ [3]) on ten scenes. The scenes include the office (1-3), the meeting room (4), the lounge (5,6), the kitchen (7), the office building (8,9), and the thicket (10). Each scene consists of three scans, except scenes 2 and 9, which are composed of two scans.
IV-B Evaluation of Point Cloud Registration Accuracy
Data. Given that the existing point cloud registration benchmark [6] lacks fiducial markers in the scenes, we construct a new test dataset, as shown in Fig. 7, with the Livox MID-40. As illustrated in the caption of Fig. 7, our test datasets cover various indoor and outdoor scenes. Indoors, multiple 16.4 cm × 16.4 cm AprilTags [11] are positioned in the environment. Outdoors, multiple 69.2 cm × 69.2 cm ArUcos [12] are placed in the environment. Note that since the LiDAR fiducial markers in this work are thin sheets of objects attached to other surfaces, they are almost invisible in the point clouds. This is infeasible if we adopt LiDARTags [19] or calibration boards [16, 17, 18] to provide fiducials.
TABLE II: comparison with LOAM methods regarding reconstruction
[…]
Livox Mapping [23] 0.0106 75.27
LOAM Livox [5] 0.0335 78.82
Traj LO [22] 0.0107 82.88
Ours 0.0030 96.22
Competitors and Metrics. Competitors include the state-of-the-art (SOTA) learning-based multiview point cloud registration method, SGHR [1], and the SOTA pairwise geometry-based method, Teaser++ [3]. We employ the root-mean-square errors (RMSEs) [32] as the metric. We also compare the inference time of the methods with an AMD Ryzen 7 5800X CPU. Results and Analysis. The qualitative and quantitative results are presented in Fig. 7 and Table I, respectively. As seen, the proposed method yields the best accuracy and efficiency. SGHR [1] successfully registers the point clouds in scenes 1 to 3 while failing in scenes 4 to 10. Its efficiency is poor on the CPU, while the inference time is around 10s to 20s on an RTX 3090 GPU. Considering that Teaser++ [3] is a pairwise method, we manually select pairs of point clouds with overlap as the inputs (the same scheme adopted in the following tests), yet it still fails in all scenes. The failure cases show SGHR’s limited generalization to unseen scenarios, and Teaser++ struggles with low overlap cases.
Figure 8: An illustration of the experimental setup and a comparison of reconstruction results of Ours, SGHR [1], and Teaser++ [3].
Figure 9: Visual comparison of the instance reconstruction. From top to bottom: ground truth, Ours, Traj LO [22], Livox Mapping [23], and LOAM Livox [5].
IV-C Evaluation of Instance Reconstruction Quality
An important application of point cloud registration is the 3D reconstruction for asset collection [33]. Thus, we evaluate instance reconstruction quality in this test. Data. The experimental setup is depicted in Fig. 8(a). Four 69.2 cm $\times$ 69.2 cm ArUco [12] markers are placed in the environment. As depicted by the yellow trajectory, the Livox MID-40 LiDAR follows a looping path to scan the vehicle, pausing at five viewpoints for a few seconds to collect relatively dense point clouds due to its sparse scanning pattern [5]. We record the rostopic of the point cloud published by the LiDAR sensor as a rosbag throughout the sampling process. Then, the same rosbag is provided to all competitors. Point cloud registration methods [1, 3], including our method, only utilize a portion of the rosbag, i.e., the input point clouds shown in Fig. 8.
TABLE III: comparison with competitors regarding localization accuracy in a degraded scene.
[…]
Traj LO [22] 0.8289 0.6077
Livox Mapping [23] 0.4854 0.0893
LOAM Livox [5] 0.8503 0.5921
SfM-M [15] 0.0551 0.0491
Ours 0.0490 0.0384
To compare the instance reconstruction quality, we use supervisely [34], a popular 3D annotation tool, to extract the point cloud of the vehicle from each scene reconstruction result. Since the manufacture and model (Mercedes-Benz GLB) of the vehicle to be reconstructed are known, a high-fidelity 3D model acquired from a 3D assets website acts as the ground truth shape. Competitors and Metrics. The competitors are point cloud registration methods (SGHR [1] and Teaser++ [3]) and LOAM methods (Traj LO[22], Livox Mapping [23], and LOAM Livox [5]). Following [35], we compute the Chamfer Distance and recall between the ground truth shape and each reconstruction result to perform a quantitative comparison of the reconstruction quality. Results and Analysis. The qualitative results for point cloud registration methods and LOAM methods are shown in Fig. 8 and Fig. 9, respectively. The quantitative results are presented in Table II. As seen in Fig. 8, SGHR is only able to align the third and fifth scans, while Teaser++ fails in all scenes even when pairs with overlap are manually selected and provided. In terms of instance reconstruction quality, as depicted in Fig. 9, our result preserves intricate details. The result of Traj LO [22] has almost no drift while the surface is noisy. The Livox Mapping [23] result exhibits noticeable drift and surface noise, while LOAM Livox [5] demonstrates severe drift. As shown in Table II, our method achieves the highest reconstruction quality.
TABLE IV: Comparison with KISS-ICP [4] regarding localization.
[…]
KISS-ICP [4] 0.1976 0.1617
Ours 0.1715 0.1394
Figure 10: Visualization of the scene reconstruction results of different methods.
Figure 11: (a): An illustration of the experimental setup. (b): Comparison of the trajectories given by different methods.
IV-D Evaluation in a Degraded Scene
As previously introduced, the adoption of fiducial markers can enhance the robustness of reconstruction and localization in degraded scenes. Therefore, in this test, we evaluate our method in a degraded scene. Data. The setup is shown in the reference of Fig. 10. This scenario has repetitive structures and weak geometric features. We attach thirteen 16.4 cm $\times$ 16.4 cm AprilTags to the wall. The LiDAR follows an arc trajectory to scan the scene, pausing at 11 viewpoints for a few seconds to collect 11 dense point clouds. We record the point cloud rostopic as a rosbag and provide it to all competitors, excluding SfM-M [15]. Point cloud registration methods only leverage a portion of the rosbag, consisting of the 11 point clouds shown in the input of Fig. 10. We also captured 72 images with an iPhone 13 to use as input for SfM-M. The ground truth trajectories are given by an OptiTrack Motion Capture system. Competitors and Metrics. Considering that all point-cloud-based competitors [1, 3, 22, 23, 5] struggle in the degraded scene and there does not exist a point-cloud-based competitor that can utilize our fiducial markers (thin sheets of letter-size paper), we add SfM-M [15], the state-of-the-art marker-based SfM method, as a competitor. This addition enables a comprehensive and meaningful comparison. We employ the RMSEs as the metric. Results and Analysis. The qualitative results and trajectory comparison are presented in Fig. 10. As depicted in Fig. 10, only our method and SfM-M [15] successfully reconstruct the scene, while the other competitors [1, 3, 22, 23, 5] fail in this degraded scene. The quantitative results shown in Table III demonstrate that the proposed approach achieves the best localization accuracy. Moreover, the performance of SfM-M is inferior to our method, which is reasonable considering LiDAR is a ranging sensor.
IV-E Evaluation of Outdoor Localization
Localization is also a crucial application of point cloud registration methods. Thus, we compare the proposed method against KISS-ICP [4] in terms of outdoor localization in this test. As shown in Fig. 11(a), Four 69.2 cm $\times$ 69.2 cm ArUco [12] markers are deployed in the environment. The vehicle, equipped with an RS-Ruby 128-beam mechanical LiDAR, follows an 8-shaped trajectory without pausing and samples 364 LiDAR scans. The ground truth is obtained from Real-Time Kinematic. As seen in Fig. 11(b) and Table IV, KISS-ICP and our method show comparable performance, with our method exhibiting less drift in the middle of the trajectory.
V Conclusion
In this work, we develop L-PR, a novel framework that exploits LiDAR fiducial markers for unordered low overlap multiview point cloud registration. L-PR consists of an adaptive threshold marker detection method and two levels of graphs. The markers in this work are thin sheets of paper attached to other surfaces, without impacting the 3D environment. The adaptive threshold method robustly detects markers even when there are dramatic changes in viewpoints among scans, thus overcoming the limitation of [20]. The proposed two-level graph efficiently aligns the low overlap multiview point clouds directly from an unordered set by optimizing a MAP problem. We conduct qualitative and quantitative experiments to demonstrate the superiorities of L-PR over competitors in terms of (1) registration accuracy, (2) instance reconstruction quality, (3) localization accuracy, and (4) robustness to the degraded scene.
Limitations and Discussions. While our proposed method shows promising results, it is constrained to scenes with deployed fiducial markers, a limitation shared by all marker-based approaches [15, 13]. Nevertheless, marker-based methods typically prioritize controllable environments, aiming to enhance reconstruction and localization in scenarios with low overlap and/or lack of texture. From this perspective, our method, developed for LiDARs, achieves the same core objective as previous methods [15, 13] designed for cameras. Moreover, as depicted in Fig. 7, the adoption of markers has no effect on the 3D geometry of the reconstructed result, given that the markers used in this work are thin sheets of paper attached to other planes.